# Frenet Immersed Finite Element Spaces on Triangular Meshes

### v1

@@ 3. Initial Construction of IFE Spaces on Triangles
${\hat{\phi}}_{0,i}^{+}(\eta,\xi)=p_{i}(\xi)+\sum\limits_{k=2}^{m}\eta^{k}\left(\sum\limits_{l=0}^{m-k}C_{k,l}^{+}p_{l}(\xi)\right)=p_{i}(\xi)+\sum_{l=1}^{l_{m}}C_{l}^{i}\mathcal{N}_{l}(\eta,\xi),~0\leq i\leq m$ (3.28)
We now use the jump conditions (2.14) to derive the formulas for computing the coefficient $C_{l}^{i},~l=1,2,\cdots,l_{m}$. These can be uniformly written as follows:
$\sum_{l=1}^{l_{m}}\left(\int_{{\hat{\Gamma}}_{K_{F}}}p_{k}(\xi)\left(\frac{\partial^{j}}{\partial\eta^{j}}\mathscr{L}\left(\mathcal{N}_{l}(0,\xi)\right)\right)d\xi\right)C_{l}^{i}=\frac{{\hat{\beta}}^{-}-{\hat{\beta}}^{+}}{{\hat{\beta}}^{+}}\int_{{\hat{\Gamma}}_{K_{F}}}p_{k}(\xi)\left(\frac{\partial^{j}}{\partial\eta^{j}}\mathscr{L}(p_{i}(\xi))\right),$ (3.29)
for $0\leq k\leq m-2-j,~0\leq j\leq m-2.$ The system (3.29) of ${l_{m}}=(m-1)m/2$ linear equations about the coefficients $\mathbf{c}^{(i)}=(C_{l}^{i})_{l=1}^{{l_{m}}}$ can be written as
@@
$A_{k,l}^{(j)}=\int_{{\hat{\Gamma}}_{K_{F}}}p_{k}(\xi)\left(\frac{\partial^{j}}{\partial\eta^{j}}\mathscr{L}\left(\mathcal{N}_{l}(0,\xi)\right)\right)d\xi,\qquad\mathbf{b}_{k}^{(j)}(i)=\int_{{\hat{\Gamma}}_{K_{F}}}p_{k}(\xi)\left(\frac{\partial^{j}}{\partial\eta^{j}}\mathscr{L}(p_{i}(\xi))\right).$
The system (3.30) is uniquely solvable, as guaranteed by Lemma 3.2. Once the remaining $m+1$ basis functions ${\hat{\phi}}_{0,i}$ are determined, together with ${\hat{\phi}}_{i,j}$ in (3.26), we obtain all $(m+1)(m+2)/2$ basis functions. Then, the Frenet IFE space on $\hat{K}_{F}$ is defined as
@@
The procedure for constructing the Frenet IFE space on $\hat{K}_{F}$ described above closely follows the construction of the Frenet-IFE space on a rectangular interface element presented in [2024AdjeridLinMeghaichi]. A key advantage of this approach is that the majority of basis functions are explicitly given in formula (3.26). Specifically, $m(m+1)/2$ of the $(m+1)(m+2)/2$ basis functions of $\hat{\mathcal{V}}^{m}_{\hat{\beta}}(\hat{K}_{F})$ are given explicitly by the simple formula (3.26), and only the remaining $m+1$ basis functions need to be constructed by solving an $l_{m}\times l_{m}$ linear system (3.29).
@@
Each Frenet–IFE function $\phi$ in the space ${\mathcal{V}}^{m}_{\beta}({K})$ defined on an interface element $K$ by (2.16) is geometrically conforming: its components $\phi^{+}$ and $\phi^{-}$ match along the entire interface curve within $K$, as illustrated in the left plot of Figure 2. In contrast, classical $P_{1}$ IFE functions [2004LiLinLinRogers] enforce continuity only at the two points where the interface intersects the element edges, as shown in the right plot of Figure 2.
Figure 2. A comparison of a $\mathbb{P}_{1}$ Frenet-IFE function (left) and a classical $\mathbb{P}_{1}$ IFE function (right)
@@ 4. General Construction for Frenet-IFE Basis on Triangles
The procedure described in the previous section for constructing a basis of $\mathcal{V}^{m}_{\beta}(K)$ uses a kind of monomial basis of $\mathbb{P}_{m}$. However, it is known that using monomial bases may lead to ill-conditioned computations, particularly when the polynomial degree becomes large. For example, using the basis $\{x^{i}\}_{i=0}^{m}$ to compute the $L^{2}$ projection onto $\mathbb{P}_{m}$ leads to a Hilbert matrix, whose condition number grows rapidly and is notoriously unfavorable for numerical approximation. In contrast, orthogonal polynomials bases, such as Legendre polynomials, often yield much better conditioned system due to intrinsic orthogonality and reduced numerical coupling among the basis functions. These considerations motivate the use of more general and computationally robust polynomial bases to construct a basis for the GC-IFE space $\mathcal{V}^{m}_{\beta}(K)$.
@@ 4.1. General Construction of GC-IFE Basis
Let $\{R_{l}\}_{l=1}^{d_{m}}$, where $d_{m}=(m+1)(m+2)/2$, be a basis of the space $\mathbb{P}_{m}$. For example, the high-degree polynomial basis with a certain orthogonal properties described in Chapter 6 of [2008HesthavenWarburton] can be used. On the Frenet reference element $\hat{K}_{F}$, we consider a set of IFE basis functions $\{\hat{\lambda}_{j}:1\leq j\leq d_{m}\}$ in the following form:
$\hat{\lambda}_{j}(\eta,\xi)=\begin{cases}\hat{\lambda}_{j}^{-}(\eta,\xi)=\displaystyle\sum\limits_{i=1}^{d_{m}}C_{ij}^{-}R_{i}(\eta,\xi),&\eta<0,\\ \hat{\lambda}_{j}^{+}(\eta,\xi)=\displaystyle\sum\limits_{i=1}^{d_{m}}C_{ij}^{+}R_{i}(\eta,\xi),&\eta>0,\end{cases}\qquad\ 0\leq j\leq d_{m}.$ (4.32)
The coefficients $C_{ij}^{\pm}$ are determined so that $\hat{\lambda}_{j}$ satisfies the interface jump conditions (2.9), (2.10), and (2.14) along ${\hat{\Gamma}}_{K_{F}}$, which give rise to the following system of equations:
$\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}R_{i}(0,\xi)\hat{p}_{k}(\xi)d\xi\right)C_{ij}^{+}=\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}R_{i}(0,\xi)\hat{p}_{k}(\xi)d\xi\right)C_{ij}^{-},\quad k=0,1,\cdots,m$ (4.33)
$\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}\partial_{\eta}R_{i}(0,\xi)\hat{p}_{k}(\xi)d\xi\right)C_{ij}^{+}=\frac{\beta^{-}}{\beta^{+}}\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}\partial_{\eta}R_{i}(0,\xi)\hat{p}_{k}(\xi)d\xi\right)C_{ij}^{-},\quad k=0,\cdots,m-1$ (4.34)
$\begin{split}\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}\frac{\partial^{n}}{\partial\eta^{n}}\mathscr{L}(R_{i}(0,\xi))\hat{p}_{k}(\xi)d\xi\right)C_{ij}^{+}=\frac{\beta^{-}}{\beta^{+}}\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}\frac{\partial^{n}}{\partial\eta^{n}}\partial_{\eta}^{n}\mathscr{L}(R_{i}(0,\xi))\hat{p}_{k}(\xi)d\xi\right)C_{ij}^{-},\\ \qquad k=0,1,\cdots,m-2-n,\quad n=0,1,\cdots m-2.\end{split}$ (4.35)
@@
where $J=\text{diag}(1,1,\cdots,1,\frac{{\hat{\beta}}^{-}}{{\hat{\beta}}^{+}},\cdots,\frac{{\hat{\beta}}^{-}}{{\hat{\beta}}^{+}})$ is a diagonal matrix whose first $m+1$ entries equal $1$ and the remaining $(d_{m}-m-1)$ entries equal ${\hat{\beta}}^{-}/{\hat{\beta}}^{+}$.
The existence and uniqueness of GC-IFE basis functions imply that, for each fixed $j$, the vector $C_{j}^{+}$ is uniquely determined from (4.36) once $C_{j}^{-}$ is given. Thus, given any basis $\hat{\lambda}_{j}^{-}:1\leq j\leq d_{m}$} of $\mathbb{P}_{m}$ on $\hat{K}_{F}^{-}$, the extension mapping (4.36) uniquely determines the corresponding functions $\hat{\lambda}_{j}^{+}:1\leq j\leq d_{m}$} on $\hat{K}_{F}^{+}$. In this way, the complete set of local GC-IFE basis functions ${\hat{\lambda}_{j}(\eta,\xi)},~1\leq j\leq d_{m}$ is generated. Furthermore, let $C^{\pm}=[C_{1}^{\pm},C_{2}^{\pm},\dots,C_{d_{m}}^{\pm}]$ denote the coefficient matrices. Then (4.36) is equivalent to the matrix identity
@@
With the basis functions constructed with this general construction procedure, we define the Frenet IFE space on $\hat{K}_{F}$ as follows:
@@
As discussed in [2025AdjeridLinMeghaichi2], the initial construction of the Frenet-based IFE basis functions $\phi_{ij}$ presented in Section 3 is a special case of this general construction. The formulation introduced here therefore not only unifies the two approaches but also provides greater flexibility in the selection of polynomial bases, which may improve conditioning and numerical performance.
@@ 4.2. Efficient generation of matrices and vectors
Next we present an approach to efficiently generate the local mass and stiffness matrices, and the source vector on a triangular interface element using GC-IFE basis functions. Let $\{\hat{\lambda}_{j}(\eta,\xi):j=1,2,\cdots,d_{m}\}$ be a basis for the reference Frenet space $\hat{\mathcal{V}}^{m}_{\hat{\beta}}(\hat{K}_{F})$ via general construction. Then $\{\lambda_{j}:=\hat{\lambda}_{j}\circ R_{\Gamma},j=1,2,\cdots,d_{m}\}$ is a basis for the local GC-IFE space ${\mathcal{V}}^{m}_{\beta}({K})$ on an interface element $K$. The local mass matrix $M_{K}=(m_{ij})$ can be written as
@@
On each piece $K^{s}$}, the integration in (4.39) is approximated by a numerical quadrature, i.e.,
@@
where $w_{k}^{s}$ and $\mathbf{x}_{k}^{s}$ are quadrature weights and nodes on $K^{s}$ and $n_{q}^{s}$ is the number of quadrature points on $K^{s}$ in a certain quadrature rule. Without loss of generality, $K^{s}$ is either a triangle or a quadrilateral, each with one curved edge; hence, associated quadrature rules available in the literature can be readily applied here.
Moreover, denote the Vandermonde matrices associated with the IFE basis $\hat{\lambda}_{i}$ and the basis $R_{i}$ of $\mathbb{P}_{m}$ associated with the nodes $\hat{\mathbf{x}}_{k}^{s}$ by
$V^{s}=\left(\hat{\lambda}_{i}(\hat{\mathbf{x}}_{k}^{s})\right)_{i=1,k=1}^{d_{m},n_{q}^{s}},\qquad L^{s}=\left(R_{i}(\hat{\mathbf{x}}_{k}^{s})\right)_{i=1,k=1}^{d_{m},n_{q}^{s}},\quad s=\pm.$ (4.41)
Then the relation (4.32) implies $V^{s}=(C^{s})^{T}L^{s}$. The mass matrix $M_{K}$ on the interface element $K$ can be approximated by $M_{K,q}$ of the following form
$M_{K}\approx M_{K,q}=\left(\sum_{s=\pm}\sum_{k=1}^{n_{q}^{s}}w_{k}^{s}\hat{\lambda}_{i}(\hat{\mathbf{x}}_{k}^{s})\hat{\lambda}_{j}(\hat{\mathbf{x}}_{k}^{s})\right)_{i,j=1}^{d_{m}}=\sum_{s=\pm}V^{s}W^{s}(V^{s})^{T}=\sum_{s=\pm}(C^{s})^{T}L^{s}W^{s}(L^{s})^{T}C^{s},$ (4.42)
@@
The subscript $q$ in $M_{K,q}$ emphasizes that the approximation depends on the quadrature rule used for integration.
@@
The procedure in (4.42) yields an efficient approximation of the local mass matrix $M_{K}$ on an interface element $K$ because the matrices $L^{s}$ are assembled by evaluating the standard polynomial basis functions $R_{i}$ at the quadrature points $\hat{\mathbf{x}}_{k}^{s}$, rather than the piecewisely defined IFE basis functions. After the coefficients $C^{\pm}$ of the GC-IFE basis functions are computed, the mass matrix $M_{K}$ is efficiently obtained by standard matrix multiplication.
This approach is applicable to the local stiffness matrix $S_{K}$ as well. In fact, the entries of $S_{K}=(S_{ij})$ can be written as
$S_{ij}=\int_{T}\beta\nabla\lambda_{i}\cdot\nabla\lambda_{j}d\mathbf{x}=\sum_{s=\pm}\int_{T^{s}}\beta^{s}\nabla(\hat{\lambda}_{i}\circ R_{\Gamma})\cdot\nabla(\hat{\lambda}_{j}\circ R_{\Gamma})d\mathbf{x}.$ (4.43)
@@
$S_{ij}$ $\approx$ $\sum_{s=\pm}\sum_{k=1}^{n_{q}^{s}}w_{k}^{s}\beta^{s}\nabla(\hat{\lambda}_{i}\circ R_{\Gamma}({\mathbf{x}}_{k}^{s}))\cdot\nabla(\hat{\lambda}_{j}\circ R_{\Gamma}({\mathbf{x}}_{k}^{s}))$ (4.44)
$=$ $\sum_{s=\pm}\sum_{k=1}^{n_{q}^{s}}w_{k}^{s}\beta^{s}\big(\hat{\lambda}_{i,\eta}(\hat{x}_{k}^{s})\hat{\lambda}_{j,\eta}(\hat{x}_{k}^{s})+\rho^{2}\hat{\lambda}_{i,\xi}(\hat{x}_{k}^{s})\hat{\lambda}_{j,\xi}(\hat{x}_{k}^{s})\big).$
@@
$V_{\eta}^{s}=\left(\hat{\lambda}_{i,\eta}(\hat{\mathbf{x}}_{k}^{s})\right)_{i=1,k=1}^{d_{m},n_{q}^{s}},\qquad V_{\xi}^{s}=\left(\hat{\lambda}_{i,\xi}(\hat{\mathbf{x}}_{k}^{s})\right)_{i=1,k=1}^{d_{m},n_{q}^{s}}$
@@
$L_{\eta}^{s}=\left(R_{i,\eta}(\hat{\mathbf{x}}_{k}^{s})\right)_{i=1,k=1}^{d_{m},n_{q}^{s}},\qquad L_{\xi}^{s}=\left(R_{i,\xi}(\hat{\mathbf{x}}_{k}^{s})\right)_{i=1,k=1}^{d_{m},n_{q}^{s}},\quad s=\pm.$

### v2

@@ 3. Initial Construction of IFE Spaces on Triangles
$\begin{split}{\hat{\phi}}_{0,i}^{+}(\eta,\xi)&=p_{i}(\xi)+\sum_{k=2}^{m}\eta^{k}\left(\sum_{l=0}^{m-k}C_{k,l}^{+}p_{l}(\xi)\right)\\ &=p_{i}(\xi)+\sum_{l=1}^{l_{m}}C_{l}^{i}\mathcal{N}_{l}(\eta,\xi),\quad 0\leq i\leq m.\end{split}$ (3.11)
We now use the jump conditions (2.10) to derive the formulas for computing the coefficient $C_{l}^{i},~l=1,2,\cdots,{\color[rgb]{0,0,0}l_{m}}$. These can be uniformly written as follows:
$\begin{split}\sum_{l=1}^{l_{m}}\left(\int_{{\hat{\Gamma}}_{K_{F}}}p_{k}(\xi)\left(\frac{\partial^{j}}{\partial\eta^{j}}\mathscr{L}\left(\mathcal{N}_{l}(0,\xi)\right)\right)d\xi\right)C_{l}^{i}\\ =\frac{{\hat{\beta}}^{-}-{\hat{\beta}}^{+}}{{\hat{\beta}}^{+}}\int_{{\hat{\Gamma}}_{K_{F}}}p_{k}(\xi)\left(\frac{\partial^{j}}{\partial\eta^{j}}\mathscr{L}(p_{i}(\xi))\right)d\xi\end{split}$ (3.12)
for $0\leq k\leq m-2-j,~0\leq j\leq m-2.$
The linear system of ${l_{m}}=(m-1)m/2$ equations (3.12) about the coefficients $C_{l}^{i},~~l=1,2,\cdots,{l_{m}}$ can be written as
@@
$A_{k,l}^{(j)}$ $=\int_{{\hat{\Gamma}}_{K_{F}}}p_{k}(\xi)\left(\frac{\partial^{j}}{\partial\eta^{j}}\mathscr{L}\left(\mathcal{N}_{l}(0,\xi)\right)\right)d\xi,$
$\mathbf{b}_{k}^{(j)}(i)$ $=\int_{{\hat{\Gamma}}_{K_{F}}}p_{k}(\xi)\left(\frac{\partial^{j}}{\partial\eta^{j}}\mathscr{L}(p_{i}(\xi))\right)d\xi.$
The system (3.13) is uniquely solvable, as guaranteed by Lemma 3.2. Once the remaining $m+1$ basis functions ${\hat{\phi}}_{0,i}$ are determined, together with ${\hat{\phi}}_{i,j}$ in (3.9), we obtain all $(m+1)(m+2)/2$ basis functions. The local space is therefore
@@
The construction developed above closely follows the construction of Frenet IFE spaces on rectangles presented in [2024AdjeridLinMeghaichi]. A key advantage of this approach is that the majority of basis functions can be constructed explicitly via (3.9). For a polynomial space of degree $m$, $m(m+1)/2$ of the $(m+1)(m+2)/2$ basis functions can be written explicitly, and only the remaining $m+1$ basis functions are obtained by solving a $l_{m}\times l_{m}$ linear system (3.12).
@@
Moreover, letting
${\hat{\phi}}_{1}(\eta,\xi)=p_{0}(\xi),~{\hat{\phi}}_{2}(\eta,\xi)=p_{1}(\xi),~{\hat{\phi}}_{3}(\eta,\xi)=\eta p_{0}(\xi)/{\hat{\beta}}(\eta,\xi),$
we can easily show that
$\hat{\mathcal{V}}^{1}_{\hat{\beta}}(\hat{K}_{F})=\text{span}\left\{{\hat{\phi}}_{1},~{\hat{\phi}}_{2},~{\hat{\phi}}_{3}\right\}.$ (3.15)
Consequently, the local $\mathbb{P}_{1}$ Frenet IFE space on an interface element $K$ has the following explicit expression, no numerical calculation is required.
${\mathcal{V}}^{1}_{\beta}({K})=\left\{{\hat{\phi}}_{1}\circ R_{\Gamma}|_{K},~{\hat{\phi}}_{2}\circ R_{\Gamma}|_{K},~{\hat{\phi}}_{3}\circ R_{\Gamma}|_{K}\right\}.$ (3.16)
This is in contrast to the $\mathbb{P}_{1}$ IFE space [2004LiLinLinRogers] whose basis functions have to numerically determined on every interface element.
Remark 3.
The Frenet–IFE shape functions $\phi$ are geometrically conforming: the pieces $\phi^{+}$ and $\phi^{-}$ match along the entire interface curve within the element, as shown in the left plot of Figure 2. In contrast, classical $\mathbb{P}_{1}$ IFE shape functions [2004LiLinLinRogers] enforce continuity only at the two intersection points where the interface cuts the element edges.
Figure 2. A comparison of Frenet-IFE shape function (left) and a classical $P_{1}$ IFE shape function (right)
@@ 4. General Construction for Frenet-IFE Basis on Triangles
It is known that monomial bases tend to produce ill-conditioned linear systems, particularly when the polynomial degree becomes large. For example, using the basis $\{x^{i}\}_{i=0}^{m}$ to compute the $L^{2}$ projection onto $\mathbb{P}_{m}$ leads to a Hilbert matrix, whose condition number grows rapidly and is notoriously unfavorable for numerical computation. In contrast, orthogonal polynomial bases, such as Legendre polynomials, often yield much better conditioned systems due to their intrinsic orthogonality and reduced numerical coupling among the basis functions. These considerations motivate the use of more general and computationally robust bases when constructing the GC-IFE space $\mathcal{V}^{m}_{\beta}(K)$.
@@ 4.1. General Construction of GC-IFE Basis
Following the construction in [2025AdjeridLinMeghaichi2], we define a set of polynomials $\{R_{l}:1\leq l\leq(m+1)(m+2)/2\}$ as follows
$R_{(t+s)(t+s+1)/2+t+1}(\eta,\xi)=q_{t}\left(\frac{\eta}{\eta_{h}}\right)p_{s}\left(\frac{\xi-\bar{\xi}}{\xi_{h}}\right),~0\leq t\leq m,0\leq s\leq m-t$ (4.1)
where $q_{0}=p_{0}\equiv 1$, and $\text{deg}(q_{t})=t$, $\text{deg}(p_{s})=s$ when $s,t\geq 1$ and such that
$q_{1}(0)=0,\quad q_{i}(0)=q_{i}^{\prime}(0)=0,~\quad 2\leq i\leq m.$
It is straightforward to verify that Let $\{R_{l}\}_{l=1}^{d_{m}}$ form a polynomial basis for the space $\mathbb{P}_{m}(K_{F})$ in the ($\eta$$\xi$)-coordinates where $d_{m}=(m+1)(m+2)/2$. We construct a set of basis functions $\{\hat{\lambda}_{j}:1\leq j\leq d_{m}\}$ on the Frenet reference element $\hat{K}_{F}$ of the following form:
$\hat{\lambda}_{j}(\eta,\xi)=\begin{cases}\hat{\lambda}_{j}^{-}(\eta,\xi)=\displaystyle\sum\limits_{i=1}^{d_{m}}C_{ij}^{-}R_{i}(\eta,\xi),&\eta<0,\\ \hat{\lambda}_{j}^{+}(\eta,\xi)=\displaystyle\sum\limits_{i=1}^{d_{m}}C_{ij}^{+}R_{i}(\eta,\xi),&\eta>0,\end{cases}\qquad\ {\color[rgb]{0,0,0}1}\leq j\leq d_{m}.$ (4.2)
The coefficients $C_{ij}^{\pm}$ are to be determined so that $\hat{\lambda}_{j}$ satisfies the interface jump conditions on ${\widehat{\Gamma}}_{K_{F}}$. These conditions give rise to the following system of equations:
$\begin{split}&\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}R_{i}(0,\xi){\color[rgb]{0,0,0}p_{k}(\xi)}\,d\xi\right)C_{ij}^{+}\\ =&\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}R_{i}(0,\xi){\color[rgb]{0,0,0}p_{k}(\xi)}\,d\xi\right)C_{ij}^{-},\quad k=0,1,\ldots,m.\end{split}$ (4.3)
$\begin{split}&\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}\partial_{\eta}R_{i}(0,\xi){\color[rgb]{0,0,0}p_{k}(\xi)}\,d\xi\right)C_{ij}^{+}\\ =&\frac{\beta^{-}}{\beta^{+}}\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}\partial_{\eta}R_{i}(0,\xi){\color[rgb]{0,0,0}p_{k}(\xi)}\,d\xi\right)C_{ij}^{-},\quad k=0,\ldots,m-1.\end{split}$ (4.4)
$\begin{split}&\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}\partial_{\eta}^{n}\mathscr{L}(R_{i}(0,\xi)){\color[rgb]{0,0,0}p_{k}(\xi)}\,d\xi\right)C_{ij}^{+}\\ =&\frac{\beta^{-}}{\beta^{+}}\sum_{i=1}^{d_{m}}\left(\int_{\xi_{0,K}}^{\xi_{1,K}}\partial_{\eta}^{n}\mathscr{L}(R_{i}(0,\xi)){\color[rgb]{0,0,0}p_{k}(\xi)}\,d\xi\right)C_{ij}^{-},\\ &\qquad\qquad k=0,\ldots,m-2-n,\quad n=0,\ldots,m-2.\end{split}$ (4.5)
@@
where $\tilde{A}$ is the coefficient matrix associated with (4.3)-(4.5), and
$J=\text{diag}(1,1,\cdots,1,\frac{\beta^{-}}{\beta^{+}},\cdots,\frac{\beta^{-}}{\beta^{+}})$
is a diagonal matrix whose first $m+1$ entries equal $1$ and the remaining $(d_{m}-m-1)$ entries equal $\beta^{-}/\beta^{+}$.
The existence and uniqueness of GC-IFE basis functions imply that, for each fixed $j$, the vector $C_{j}^{+}$ is uniquely determined from (4.6) once $C_{j}^{-}$ is given. Thus, given any basis $\hat{\lambda}_{j}^{-}:1\leq j\leq d_{m}$} of $\mathbb{P}_{m}$ on $\hat{K}_{F}^{-}$, the extension mapping (4.6) uniquely determines the corresponding functions $\hat{\lambda}_{j}^{+}:1\leq j\leq d_{m}$} on $\hat{K}_{F}^{+}$. In this way, the complete set of local GC-IFE basis functions $\hat{\lambda}_{j}(\eta,\xi)$}, $1\leq j\leq d_{m}$ is generated.
Let $C^{\pm}=[C_{1}^{\pm},C_{2}^{\pm},\dots,C_{d_{m}}^{\pm}]$ denote the coefficient matrices. Then (4.6) is equivalent to the matrix identity
@@
Using the general construction, the local space can be written as
@@
As discussed in [2025AdjeridLinMeghaichi2], the initial construction of the Frenet-based IFE basis functions $\phi_{ij}$ presented in Section 3 arises as a special case of this more general GC-IFE framework. The formulation introduced here therefore not only unifies the two constructions but also provides greater flexibility in selecting polynomial bases that may improve conditioning and numerical performance.
@@ 4.2. Efficient generation of matrices and vectors
Next we present an approach to efficiently generate local mass and stiffness matrices and source vectors on a triangular interface element using GC-IFE basis functions. Let $\{\hat{\lambda}_{j}(\eta,\xi):j=1,2,\cdots,d_{m}\}$ be a basis for the reference Frenet space $\hat{\mathcal{V}}^{m}_{\hat{\beta}}(\hat{K}_{F})$ via general construction. Then $\{\lambda_{j}:=\hat{\lambda}_{j}\circ R_{\Gamma},j=1,2,\cdots,d_{m}\}$ is a basis for the local GC-IFE space $V^{m}(K)$ on an interface element $K$. The local mass matrix $M_{K}=(m_{ij})$ can be written as
@@
On each piece $K^{s}$}, the integration (4.9) is approximated by numerical quadratures, i.e.,
@@
where $w_{k}^{s}$ and $\mathbf{x}_{k}^{s}$ are quadrature weights and nodes on $K^{s}$ and $n_{q}^{s}$ is the number of quadrature points on $K^{s}$ in a certain quadrature rule. Note that $K^{s}$ is either a curved triangle or quadrilateral. The Stroud quadrature rule introduced in Section 3.2 of [2025AdjeridLinMeghaichi2] for rectangular meshes is also applicable in this case.
Denote the Vandermonde matrices associated with the basis $\hat{\lambda}_{i}$ and the basis $R_{i}$ of $\mathbb{P}_{m}$ to be
$V^{s}=\left(\hat{\lambda}_{i}(\hat{\mathbf{x}}_{k}^{s})\right)_{k=1,i=1}^{n_{q}^{s},d_{m}},\qquad L^{s}=\left(R_{i}(\hat{\mathbf{x}}_{k}^{s})\right)_{k=1,i=1}^{n_{q}^{s},d_{m}},\quad s=\pm.$ (4.11)
Then the relation (4.2) implies $V^{s}=L^{s}C^{s}$. The mass matrix $M_{K}$ on the interface element $K$ can be approximated by $M_{K,q}$ of the following form
$\begin{split}{\color[rgb]{0,0,0}M_{K}\approx M_{K,q}}&=\left(\sum_{s=\pm}\sum_{k=1}^{n_{q}^{s}}w_{k}^{s}\hat{\lambda}_{i}(\hat{\mathbf{x}}_{k}^{s})\hat{\lambda}_{j}(\hat{\mathbf{x}}_{k}^{s})\right)_{i,j=1}^{d_{m}}\\ &=\sum_{s=\pm}(V^{s})^{T}W^{s}V^{s}=\sum_{s=\pm}(C^{s})^{T}(L^{s})^{T}W^{s}L^{s}C^{s}.\end{split}$ (4.12)
@@
The notation $M_{K,q}$ emphasizes that the approximation depends on the quadrature rule used for integration.
@@
The procedure in (4.12) yields an efficient approximation of the element mass matrix $M_{K}$, because the matrices $L^{s}$ are assembled by evaluating the standard polynomial basis functions $R_{i}$ at the quadrature points $\hat{\mathbf{x}}_{k}^{s}$, rather than using piecewise polynomials. After the coefficients of the GC-IFE shape functions $C^{\pm}$ are computed, the mass matrix $M_{K}$ is obtained by standard matrix multiplication.
The construction of the local right hand side vector $\mathbf{f}_{K}$ on an interface element follows a similar procedure. In fact,
${\bf f}_{K}\approx{\bf f}_{K,q}=\sum_{s=\pm}(V^{s})^{T}W^{s}\mathbf{r}^{s}=\sum_{s=\pm}(C^{s})^{T}(L^{s})^{T}W^{s}\mathbf{r}^{s},\qquad$ (4.13)
where
$\mathbf{r}^{s}=\Big(f(\mathbf{x}_{1}^{s}),f(\mathbf{x}_{2}^{s}),\cdots,f(\mathbf{x}_{n_{q}^{s}}^{s})\Big)^{T}.$
This approach is applicable to the stiffness matrix $S$ as well. The entries of ${\color[rgb]{0,0,0}S_{K}}=(S_{ij})$ can be written as
$S_{ij}=\int_{K}\beta\nabla\lambda_{i}\cdot\nabla\lambda_{j}d\mathbf{x}=\sum_{s=\pm}\int_{K^{s}}\beta^{s}\nabla(\hat{\lambda}_{i}\circ R_{\Gamma})\cdot\nabla(\hat{\lambda}_{j}\circ R_{\Gamma})d\mathbf{x}.$ (4.14)
@@
$\begin{split}S_{ij}&\approx\sum_{s=\pm}\sum_{k=1}^{n_{q}^{s}}w_{k}^{s}\beta^{s}\nabla(\hat{\lambda}_{i}\circ R_{\Gamma}(\mathbf{x}_{k}^{s}))\cdot\nabla(\hat{\lambda}_{j}\circ R_{\Gamma}(\mathbf{x}_{k}^{s}))\\ &=\sum_{s=\pm}\sum_{k=1}^{n_{q}^{s}}w_{k}^{s}\beta^{s}\big(\hat{\lambda}_{i,\eta}(\hat{x}_{k}^{s})\hat{\lambda}_{j,\eta}(\hat{x}_{k}^{s})+\rho^{2}\hat{\lambda}_{i,\xi}(\hat{x}_{k}^{s})\hat{\lambda}_{j,\xi}(\hat{x}_{k}^{s})\big).\end{split}$ (4.15)
@@
$V_{\eta}^{s}=\left(\hat{\lambda}_{i,\eta}(\hat{\mathbf{x}}_{k}^{s})\right)_{k=1,i=1}^{n_{q}^{s},d_{m}},\qquad V_{\xi}^{s}=\left(\hat{\lambda}_{i,\xi}(\hat{\mathbf{x}}_{k}^{s})\right)_{k=1,i=1}^{n_{q}^{s},d_{m}}$
@@
$L_{\eta}^{s}=\left(R_{i,\eta}(\hat{\mathbf{x}}_{k}^{s})\right)_{k=1,i=1}^{n_{q}^{s},d_{m}},\qquad L_{\xi}^{s}=\left(R_{i,\xi}(\hat{\mathbf{x}}_{k}^{s})\right)_{k=1,i=1}^{n_{q}^{s},d_{m}},\quad s=\pm.$